OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1).

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

No, that's why I'm asking these questions.

OpenStudy (anonymous):

You can download one from the internet. Easy peasy.

OpenStudy (anonymous):

I'm not using my own computer right now.

OpenStudy (anonymous):

first we have to find our slope, m m= -1-1/5-4=-2

OpenStudy (anonymous):

next we substitute m into our slope intercept equation y=-2x+b and plug in one of the given coordinates (5,-1) -1=-2(5)+b -1=-10+b b=9 therefore our slope intercept equation is y=-2x+9 :]

OpenStudy (anonymous):

Thanks so much love :]